OpenStudy (elenathehomeschooler):

Help please! Find an algebraic expression equivalent to the following expressions: cos(tan^-1 x) cot(arcsin x)

OpenStudy (elenathehomeschooler):

@jim_thompson5910 can you help ?

OpenStudy (jdoe0001):

hmmm, do you recall the ranges for inverse trig functios? that is \(\textit{Inverse Trigonometric Identities} \\ \quad \\ \begin{array}{cccl} Function&{\color{brown}{ Domain}}&{\color{blue}{ Range}}\\ \hline\\ {\color{blue}{ y}}=sin^{-1}({\color{brown}{ \theta}})&-1\le {\color{brown}{ \theta}} \le 1&-\frac{\pi}{2}\le {\color{blue}{ y}}\le \frac{\pi}{2} \\ \quad \\ {\color{blue}{ y}}=cos^{-1}({\color{brown}{ \theta}})&-1\le {\color{brown}{ \theta}} \le 1& 0 \le {\color{blue}{ y}}\le \pi \\ \quad \\ {\color{blue}{ y}}=tan^{-1}({\color{brown}{ \theta}})&-\infty\le {\color{brown}{ \theta}} \le +\infty &-\frac{\pi}{2}\le {\color{blue}{ y}}\le \frac{\pi}{2} \end{array}\)

OpenStudy (elenathehomeschooler):

yes

OpenStudy (anonymous):

\[put~\tan^{-1} x=y,x=\tan y\]